# SemHARQ: Semantic-Aware HARQ for Multi-task Semantic Communications

### v1

@@ V-B Simulation settings and numerical analysis
The detailed network settings of the proposed SemHARQ are shown in Table I. The system is evaluated under several SNR levels and the average power budget P is set to $1$. During training, the input images are first resized to $256\times 256$. DenseNet-121 is utilized as our backbone deep neural network (DNN), which is optimized by Adam [40]. $\lambda_{r}$, $\lambda_{c}$, $\lambda_{t}$ are set to $1,0.125,0.125$, respectively. The training batch and testing batch are set to $16$ and $100$, respectively. Besides, the learning rate is initialized as 1e44-4- 4, and learning rate decay is applied. By adjusting the distortion threshold $t$ in III-A, the number of features to be retransmitted is approximately the same as those for incremental transmission on average. In addition, the given conversion ratio $\theta_{0}$ in IV-A is set to $1$ dB, and the given maximum SNR $\theta_{\text{SNR}}$ can be determined by $\theta_{\mathrm{SNR}}=\theta_{0}\left[1+\log{M}\right]$, where $\log{M}$ is the maximum entropy, $M$ denotes the total number of categories.
@@
Note that the IK-S-HARQ and RT-S-HARQ are both DL-based and JSCC is applied to achieve the best results.
Fig. 8 presents the vehicle ReID performance w.r.t. different methods. It is obvious that the semantic-aware HARQ methods outperform the traditional method in both rank-1 accuracy and mAP, which verifies the superiority of SemCom in transmitting the meaning of information efficiently. Furthermore, the proposed SemHARQ achieves the best performance among three semantic-aware HARQ methods, followed by IK-S-HARQ and RT-S-HARQ. The rank-1 accuracy gaps between SemHARQ and IK-S-HARQ and RT-S-HARQ at $-2$ dB are $18.50\%$ and $30.97\%$, respectively, while in terms of the mAP, the gaps are $21.78\%$ and $43.61\%$, respectively. The gain is attributed to the precise evaluation of the feature distortion in SemHARQ, where only distorted features are allowed to be retransmitted, improving the efficiency of channel resource utilization significantly. Note that the percentage of mAP is generally lower than the rank-1 accuracy. For example, the rank-1 accuracy of SemHARQ is $68.22\%$, yet the corresponding mAP is only $42.94\%$ at $-2$ dB. The reason is that mAP combines accuracy and recall for different categories of the recognition task, which is usually more rigorous than rank-1 accuracy.
@@
At the same time, the accuracy of the other two classification tasks is illustrated in Fig. 9. The performance of SemHARQ is superior to other methods, especially at low SNR regime. Specifically, at $-6$ dB, in terms of color classification, SemHARQ improves the performance by $10.28\%$, $40.88\%$ and $166.48\%$, compared with the IK-S-HARQ, the RT-S-HARQ and the traditional HARQ, respectively. For type classification, the performance is improved by $8.83\%$, $15.46\%$, and $73.01\%$, respectively. Note that the performance gain progressively reduces as SNR increases, which results from the correct reception of most of the information under good channel conditions. On this basis, the retransmissions become less meaningful.
To further verify the effectiveness of the FDE network in SemHARQ, we conduct an ablation study for comparison. The distorted features are directly identified from the received feature vector ${\bm{\hat{z}}}_{j}$ using multi-dimension Gumbel-softmax in the comparison method. Accordingly, $\mathcal{F}$ and mutual information estimation discussed in III-A are not involved, and FDE network training is not taken into account. As shown in Fig. 10, the multi-task accuracy w.r.t. different channel conditions is demonstrated. It shows that our proposed SemHARQ achieves better task performance, especially with poor SNRs. Specifically, at $-6$ dB, the multi-task performance has been improved by $67.52\%$, $2.69\%$, and $13.01\%$ by using the FDE network, respectively. The reason is that the FDE network enhances the distortion evaluation ability of the system by measuring the contribution of the received semantics to the mutual information at each transmission. The obtained distortion levels quantify the degree of feature distortion, providing precise guidance on retransmission. Therefore, the efficiency of retransmission is improved dramatically.
@@
TABLE II: Performance comparison of the proposed FIR method and the SST method.
[…]
Rank-1 acc. Color acc. Type acc. Rank-1 acc. Color acc. Type acc.
$-6$ 16.97 88.95 76.68 1.61 42.54 47.51
$-4$ 19.65 90.81 83.15 5.96 70.76 70.54
$-2$ 71.22 91.30 88.81 27.83 82.24 82.15
00 79.94 92.17 91.07 68.00 90.72 87.67
$2$ 85.74 93.75 91.64 83.91 92.67 90.97
$4$ 89.94 93.75 91.75 88.97 92.80 91.45
$6$ 90.01 94.21 91.85 89.52 93.47 91.29
$8$ 90.94 94.55 91.92 90.01 93.54 91.54
The impact of FIR is also presented in Table II. The task performance is compared with the sequentially selected transmission (SST) method which transmits the first $B$ features without considering the importance levels at each transmission. Both FIR and SST perform scalable transmission w.r.t. various SNR levels. It is obvious that the FIR method surpasses the SST method significantly for all three tasks. Particularly, at $-2$ dB, the recognition accuracy increases by $155.91\%$, $11.02\%$, and $8.11\%$ in the three tasks, respectively. Note that vehicle ReID is more challenging due to the high intra-class variability coupled with small inter-class variability. However, greater advancements are made on the vehicle ReID than the remaining two tasks, demonstrating the superiority of the proposed system. One of the reasons is that the proposed FIR can build the relationship between the importance distribution of the semantic information and the multi-task predictions, thus effectively finding the optimal feature selection policy for partial semantics transmissions.
Besides, to validate the effectiveness of the proposed SemHARQ more comprehensively, the gain of the system is investigated by increasing the number of retransmissions progressively. Fig. 11 presents the impact of retransmissions on the task performance in the case of $\text{SNR}=-6$ dB. Note that as the number of retransmissions increases, the system performance improves obviously. Particularly, the performance of vehicle ReID after three retransmissions is almost triple than that of the no-retransmission method, which further proves the validity of SemHARQ.
Moreover, to confirm the effectiveness of the retransmission identification module, the average numbers of retransmissions are summarized in Fig. 12 w.r.t. multiple SNR levels, where the maximum number is set large enough. As shown in Fig. 12, the retransmissions are more frequent under harsh channel conditions to meet the criteria in (16). The average number of retransmissions at $-6$ dB is almost 18 times higher than at $8$ dB when $\theta_{0}$, the given conversion ratio between the uncertainty and SNR, is set to $1$ dB. Meanwhile, the need for retransmissions reduces when $\theta_{0}$ becomes lower accordingly, which represents that the receiver’s demand for necessary information reception has been decreased.

### v2

@@ V-B Simulation settings and numerical analysis
The detailed network settings of the proposed SemHARQ are shown in Table I. The system is evaluated under several SNR levels and the average power budget P is set to $1$. DenseNet-121 is utilized as our backbone deep neural network (DNN), which is optimized by Adam [43]. $\lambda_{r}$, $\lambda_{c}$, $\lambda_{t}$ are set to $1,0.125,0.125$, respectively. For a given SNR, the entire training process consists of 135 epochs, including 30 epochs for pre-training the FDE network to estimate feature distortion, followed by 40 epochs for training the whole SemHARQ system without scalable feature selection to obtain feature importance scores, and another 65 epochs for retraining the SemHARQ model with scalable feature selection based on the learned importance scores. The training batch and testing batch are set to 32 and 100, respectively. Both training and testing are conducted on a single NVIDIA V100 GPU. The entire model is trained using a dataset of approximately 40,000 RGB images with a resolution of $3\times 256\times 256$. To further illustrate the convergence behavior of our model, a training loss curve of SemHARQ training with SNR = 0 dB is shown in Fig. 8. It can be observed that the training loss decreases smoothly and steadily during the three training phases as discussed in Algorithm 1, eventually converging around epoch 25, epoch 60 and epoch 130, respectively. Meanwhile, around 3.6 GB of memory is consumed during testing on images with a resolution of $3\times 256\times 256$. Note that techniques such as model pruning [51] and knowledge distillation [52] can be adopted to further reduce memory consumption in practical applications. Besides, the learning rate is initialized as 1e4-4, and learning rate decay is applied. In addition, the conversion ratio between the uncertainty and the receive-SNR of the combined information $\theta_{0}$ in Eq. (19) is set to $1$ dB, and the maximum receive-SNR $\theta_{\text{SNR}}$ can be determined by $\theta_{\mathrm{SNR}}=\theta_{0}\left[1+\log{M}\right]$, where $\log{M}$ is the maximum entropy, with $M$ denoting the total number of categories. Moreover, to trade off the robustness and the diversity of the received semantics, the number of features to be retransmitted is approximately the same as those for incremental transmission on average by adjusting the distortion threshold $t$ in Eq. (11). Note that the value of $t$ can also be readily changed. Particularly, to explore the impact of $t$, the multi-task performance of different retransmission ratios after adjusting $t$ is shown in Table II, where the retransmission ratio represents the average ratio of the retransmitted features to all the received features. It is obvious that better performance is achieved when the ratio is set to $0.5$. The reason is that the semantic error correction and new information transmission are better traded off, enhancing the communication efficiency.
Figure 8: The loss curve of SemHARQ training with SNR = 0 dB.
TABLE II: Multi-task accuracy of different retransmission ratios.
Retransmission ratio
SNR [dB] 0.25 0.5 0.75
Rank-1 Color Type Rank-1 Color Type Rank-1 Color Type
6-6 17.61 88.09 76.48 20.97 88.95 80.68 18.85 88.10 78.86
4-4 19.45 89.69 84.30 23.65 90.81 85.15 20.89 89.75 85.26
2-2 60.98 90.49 88.47 68.22 91.30 88.81 63.35 90.75 88.55
0 72.94 91.17 90.07 76.94 92.17 91.07 73.69 92.02 90.84
22 85.11 93.19 91.79 85.74 93.75 91.64 85.00 92.25 91.53
44 89.78 93.51 91.92 89.94 93.78 91.75 89.90 93.06 91.80
66 90.40 93.61 92.36 90.58 94.21 91.85 90.41 93.40 92.32
88 90.41 93.72 92.23 91.54 94.15 92.52 91.03 93.80 92.55
@@
Note that the comparison only focuses on the HARQ method at the semantic feature level, which is independent of the source modality. The proposed SemHARQ and the semantic-aware HARQ baselines are all DL-based with the same backbone DNN, i.e., DenseNet-121. Meanwhile, the JSCC and scalable feature selector with the same structures are applied to achieve the best results.
Fig. 9 presents the vehicle ReID performance w.r.t. different methods. It is obvious that the semantic-aware HARQ methods outperform the traditional method in both rank-1 accuracy and mAP, which verifies the superiority of SemCom in transmitting the meaning of information efficiently. Furthermore, the proposed SemHARQ achieves the best performance among the three semantic-aware HARQ methods, followed by IK-S-HARQ and RT-S-HARQ. The rank-1 accuracy gaps between SemHARQ and IK-S-HARQ and RT-S-HARQ at 6-6 dB are $49.79\%$ and $77.71\%$, respectively, while in terms of the mAP, the gaps are $67.93\%$ and $139.17\%$, respectively. The gain is attributed to the precise evaluation of the feature distortion in SemHARQ, where only distorted features are allowed to be retransmitted, improving the efficiency of channel resource utilization significantly, especially for the low SNR regimes.
To further demonstrate the performance of SemHARQ at any given SNR regime, the task performance of the perfect transmission is shown in Fig. 9 and Fig. 10, which are obtained by performing the tasks directly using the extracted semantic feature vector without feature selection and transmission (i.e., transmission conditions are good enough to ensure perfect reconstruction of transmitted symbols at the receiver, as shown in the highest line in Fig. 9 and Fig. 10). Note that to ensure a fair comparison, the same backbone DNN as SemHARQ is utilized. At low SNR regime, the transmission capacity is severely constrained and the transmitted features are heavily corrupted by the channel fading and noises. Therefore, under the constraint of the maximum number of retransmissions, the performance of SemHARQ is relatively poor. For example, at -6 dB, the SemHARQ exhibits a $70.97\%$ absolute reduction in rank-1 accuracy compared to the perfect transmission method in Fig. 9. However, this limitation is inherent to all HARQ methods operating under such transmission conditions. Particularly, the superiority of our SemHARQ lies in combining the retransmissions of distorted features with the incremental transmissions, which significantly improves the task performance compared to other HARQ methods, despite the capacity constraint. It is obvious that at high SNR regime, the performance of SemHARQ is very close to the perfect transmission, which results from the correct reception of most of the information under good channel conditions and high transmission capacity. For example, as presented in Fig. 9, the rank-1 accuracy of the SemHARQ only reduces $0.44\%$ compared to the perfect transmission when SNR is $8$ dB. Furthermore, note that the performance gains of SemHARQ over IK-S-HARQ and RT-S-HARQ are also marginal when SNR is high. The reason is that the decoding is mostly successful after the first transmission, and thus, retransmission seldomly triggers under this condition. As a result, the superiority of SemHARQ over other semantic-aware HARQ methods degrades.
@@
At the same time, the accuracy of the other two classification tasks is illustrated in Fig. 10. The performance of SemHARQ is superior to other methods, especially for the low SNR regimes. Specifically, at 6-6 dB, in terms of color classification, SemHARQ improves the performance by $10.28\%$, $40.88\%$, and $166.48\%$, compared with the IK-S-HARQ, the RT-S-HARQ and the traditional HARQ, respectively. For type classification, the performance is improved by $14.50\%$, $33.55\%$, and $82.04\%$, respectively. Note that since the classification task is less difficult than the ReID task, the accuracy is not affected by a certain amount of semantic errors. Therefore, the semantic baselines perform well even without feature distortion evaluation, leading to less performance gain. Additionally, the performance gain progressively reduces as SNR increases due to the correct decoding of the received semantics under good channel conditions. On this basis, the retransmissions become less meaningful.
TABLE III: Performance comparison of the proposed FIR method and the SST method.
SNR [dB] Proposed FIR ($\%$) SST ($\%$)
Rank-1 acc. Color acc. Type acc. Rank-1 acc. Color acc. Type acc.
6-6 20.97 88.95 80.68 1.61 42.54 47.51
4-4 23.65 90.81 85.15 5.96 70.76 70.54
2-2 68.22 91.30 88.81 27.83 82.24 82.15
0 76.94 92.17 91.07 68.00 90.72 87.67
22 85.74 93.75 91.64 83.91 92.67 90.97
44 89.94 93.78 91.75 88.97 92.80 91.45
66 90.58 94.21 91.85 89.52 93.47 91.29
88 91.54 94.15 92.52 90.01 93.54 91.54
To further verify the effectiveness of the FDE network in SemHARQ, we conduct an ablation study for comparison. The distorted features are directly identified from the received feature vector ${\bm{\hat{z}}}^{j}$ using multi-dimension Gumbel-softmax in the comparison method. Accordingly, $\mathcal{F}$ and mutual information estimation discussed in III-A are not involved, and FDE network training is not taken into account. As shown in Fig. 11, the multi-task accuracy w.r.t. different channel conditions is demonstrated. It shows that our proposed SemHARQ achieves better task performance, especially with poor SNRs. Specifically, at 6-6 dB, the multi-task performance has been improved by $107.00\%$, $2.69\%$, and $18.91\%$ by using the FDE network, respectively. The reason is that the FDE network enhances the distortion evaluation ability of the system by measuring the contribution of the received semantics to the mutual information at each transmission. The obtained distortion levels quantify the degree of feature distortion, providing precise guidance on retransmission. Therefore, the efficiency of retransmission is improved dramatically.
@@
The impact of FIR is also presented in Table III. The task performance is compared with the sequentially selected transmission (SST) method which transmits the first $B$ features without considering the importance levels at each transmission. Both FIR and SST perform scalable transmission w.r.t. various SNR levels. It is obvious that the FIR method surpasses the SST method significantly for all three tasks. Particularly, at 2-2 dB, the recognition accuracy increases by $45.13\%, 11$$1.02\%$, and $8.11\%$ in the three tasks, respectively. Note that vehicle ReID is more challenging due to the high intra-class variability coupled with small inter-class variability. However, greater advancements are made on the vehicle ReID than the remaining two tasks, demonstrating the superiority of the proposed system. One of the reasons is that the proposed FIR can build the relationship between the importance distribution of the semantic information and the multi-task predictions, thus effectively finding the optimal feature selection policy for partial semantics transmissions.
Besides, to validate the effectiveness of the proposed SemHARQ more comprehensively, the gain of the system is investigated by increasing the number of retransmissions progressively. Fig. 12 presents the impact of retransmissions on the task performance in the case of $\text{SNR}=-6$ dB. Note that as the number of retransmissions increases, the system performance improves obviously. Particularly, the performance of vehicle ReID after three retransmissions is almost four times that of the no-retransmission method, which further proves the validity of SemHARQ.
Moreover, to confirm the effectiveness of the retransmission identification module, the average numbers of retransmissions are summarized in Fig. 13 w.r.t. multiple SNR levels, where the maximum number is set large enough. As shown in Fig. 13, the retransmissions are more frequent under harsh channel conditions to meet the criteria in Eq. (19). The average number of retransmissions at 6-6 dB is almost 18 times higher than at $8$ dB in SemHARQ when $\theta_{0}$, the given conversion ratio between the uncertainty and SNR, is set to $1$ dB. Meanwhile, the need for retransmissions reduces when $\theta_{0}$ becomes lower accordingly, which represents that the receiver’s demand for necessary information reception has been decreased. Furthermore, note that the RT-S-HARQ and IK-S-HARQ methods require more frequent retransmissions than SemHARQ. Particularly, compared to SemHARQ, the numbers of retransmissions for RT-S-HARQ and IK-S-HARQ increase by $19.82\%$ and $12.95\%$ when SNR is set to 6-6 dB, respectively. The reason is that SemHARQ has higher resource utilization when combining the retransmission of distorted features and the incremental transmission of new features, reducing the need for retransmission.
@@
TABLE IV: Computational complexity and number of parameters of each module in SemHARQ.
Transmission system Complexity Parameters
Multi-task semantic encoder $O(W\cdot H\cdot N)$ $6.95~\text{M}$
JSC-Encoder $O(N^{2})$ $2.10~\text{M}$
JSC-Decoder $O(J\cdot B\cdot N)$ $17.83~\text{M}$
Semantic decoders and task performers $O\big{(}N\cdot(M_{1}+M_{2}+M_{3})\big{)}$ $3.68~\text{M}$
FDE network $O(N^{2})$ $1.49~\text{M}$
TABLE V: Execution delay of each module in SemHARQ with batch size 100.
Transmission system Execution delay
Multi-task semantic encoder $55.27~\text{ms}$
JSC-Encoder $0.13~\text{ms}$
Retransmissions $2.2~\text{ms}$
JSC-Decoder $0.14~\text{ms}$
Semantic decoders and task performers $0.27~\text{ms}$
FDE network $0.16~\text{ms}$
Overall inference latency $\color[rgb]{0,0,0}58.17~\text{ms}$}
V-C Discussions
V-C1 Computational and implementation complexity
The computational complexity and number of parameters of different modules of SemHARQ are shown in Table IV. Specifically, the complexity of the multi-task semantic encoder is mainly impacted by the input and output dimensions of the DNN, where $W$, $H$ are the width and height of the input images, respectively, and $N$ denotes the size of the encoded semantic feature vector [14]. Compared with the JSC-Encoder, whose complexity is only dependent on the length of the semantic vector, the JSC-Decoder becomes more complicated as the total number of retransmissions $J$ increases. The reason is that the received feature vectors after all the transmissions are concatenated and decoded to the original length $N$ by the JSC-Decoder. Note that $B$ represents the dimension of the transmitted vector after scalable feature selection. Moreover, the complexity of semantic decoders and task performers is given by $O\big{(}N\cdot(M_{1}+M_{2}+M_{3})\big{)}$ [32], where $M_{1},M_{2},$ and $M_{3}$ indicate the number of categories of the considered three tasks, respectively. It is obvious that more complex tasks require more complicated semantic decoders and task performers. Finally, the FDE network consists of multiple FC layers, whose complexity is determined by the input dimension $N$.
Furthermore, as shown in Table IV, the JSC-Decoder has the largest number of parameters due to the multiple transmissions. Meanwhile, since the DNN is used to extract semantic features, the parameter count of the multi-task semantic encoder is $6.95$ M, approximately one-third of that of the JSC-Decoder. Other modules have fewer parameters because they only consist of FC layers with low input dimensions. Note that compared with the semantic-aware HARQ baselines, the only complexity increase of our proposed SemHARQ lies in the FDE network, which is negligible and acceptable.
Finally, an evaluation of the execution delay is conducted for different modules in the proposed SemHARQ system in Table V. The delay is measured in terms of average time taken for each module to process a batch of data at an NVIDIA V100 GPU, where the batch size is set to 100. As shown in Table V, the execution delay of the multi-task semantic encoder is the highest, approximately 300 times that of the other modules. The reason is that the multi-task semantic encoder is composed of DenseNet-121, which has 121 layers of convolutional neural network, resulting in relatively large computational cost. However, the other modules like JSC-Encoder/Decoder, semantic decoders, task performers, each consists of two fully connected layers, respectively, while the FDE network comprises three, which is very lightweight with low execution delay. Additionally, the average retransmission delay under different SNRs is $2.2~\text{ms}$ with batch size 100, which is acceptable in practical systems.
TABLE VI: Classification accuracy of different semantic-aware HARQ methods on ImageNet.
SNR [dB] SemHARQ (Ours) IK-S-HARQ RT-S-HARQ
6-6 21.36 15.85 12.16
2-2 69.22 58.32 54.68
22 84.53 83.11 78.54
66 91.56 90.98 90.36
Figure 14: The mutual information of the upper bound and the proposed SemHARQ w.r.t. different SNRs.
V-C2 FDE module
Note that the FDE network - based on MINE - is the key module in the proposed SemHARQ. As analyzed in [53, 54, 40], when dealing with noisy variables with low relevance, MINE may suffer from high variance, particularly when handling high-dimensional inputs with insufficient batch size of training samples. For example, for noisy vectors manifesting Gaussian distributions, variance in the estimation tends to increase as the true mutual information grows when the dimension of the data is about 100 with batch size less than 64 [54, 53]. However, the task-related semantic features in SemCom frameworks exhibit higher semantic relevance and less randomness compared to randomly generated variables. Therefore, the uncertainty in the mutual information estimation is significantly reduced, leading to a stable MINE module, as in [55]. Note that the major variance introduced in SemHARQ lies in the channel fading and additive noise. To mitigate the impact, the transmission loss over the channel is combined with the mutual information loss during the training of the MINE module, which improves the scalability of different channel states. Additionally, the proposed FDE network is task-specific in our framework, which is consistent with the existing task-oriented semantic communication methods [9, 12]. However, the FDE methodology is general and can be applied to new semantic tasks by retraining or transfer learning techniques [56]. For example, the pretrained FDE network can be efficiently adapted to new tasks via fine-tuning or domain adaptation using a few new labels [56].
To further verify the scalability of the proposed SemHARQ on more complicated scenarios, the classification accuracy of different semantic-aware HARQ methods on the ImageNet dataset with rician fading channel is shown in Table VI. Note that ImageNet is one of the most comprehensive image datasets, encompassing a larger number of categories than our IoV scenario, with 1,000 diverse categories and varying data distributions [57]. As presented in Table VI, our proposed SemHARQ achieves the best performance among three semantic-aware HARQ methods, followed by IK-S-HARQ and RT-S-HARQ, especially for the low SNR regimes. Specifically, the classification accuracy gaps between SemHARQ and IK-S-HARQ and RT-S-HARQ at 6-6 dB are $34.76\%$ and $75.66\%$, respectively. The performance gain is attributed to SemHARQ’s design in accurately evaluating the incurring feature distortion because of an accurate estimation by MINE, demonstrating the practicality of MINE for semantic features on diverse datasets.
V-C3 Bounds of SemHARQ
To bound the performance margin between the optimal scheme and the proposed SemHARQ, we first depict the performance upper bound of the optimal scheme using mutual information. Given the optimal method that noisy transmission is not considered, the maximum mutual information between the semantic encoded feature vector and the multi-task prediction vectors is defined as follows:
${\color[rgb]{0,0,0}I_{\text{DL}}=\sum_{k=1}^{K}\max_{\alpha,\beta,\alpha^{-1},{\beta_{k}^{-1}},{\phi_{k}}}I\left(\bm{s}^{\prime};\bm{c}^{\prime}_{k}\right),}$ (30)
where $K$ represents the number of tasks, $\alpha,\beta,\alpha^{-1},{\beta_{k}^{-1}},{\phi_{k}}$ are the parameter sets of the semantic encoder, JSC-Encoder, JSC-Decoder, and the semantic decoder and task performer for the $k$-th task, respectively. Particularly, $I\left(\bm{s}^{\prime};\bm{c}^{\prime}_{k}\right)$ denotes the mutual information between the semantic encoded feature vector $\bm{s}^{\prime}$ and the $k$-th task prediction vector $\bm{c}^{\prime}_{k}$. $I_{\text{DL}}$ indicates the optimal performance of the adopted deep learning model from the perspective of information theory. On this basis, we can obtain the ideal mutual information achievable at any given SNR, where Rician channel model is considered. Note that to ensure a fair comparison with our proposed SemHARQ, retransmission scheme is also utilized when calculating the theoretical limit of the mutual information under different SNRs. Specifically, the upper bound of the mutual information between the transmitter and the receiver is the total channel capacity with multiple transmissions, which can be represented as follows,
$I_{\text{U}}$ $=\min(\frac{I_{\text{DL}}}{C},G_{\text{max}})\times C=\min({I_{\text{DL}}},G_{\text{max}}\times C),$ (31)
where $C=\log_{2}\left(1+\text{SNR}\cdot\left\|\bm{h}\right\|^{2}\right)$ is the channel capacity with the fading coefficient vector $\bm{h}$, $G_{\text{max}}$ is the considered maximum number of transmissions, and $\min(\frac{I_{\text{DL}}}{C},G_{\text{max}})$ denotes the actual number of transmissions. This indicates that $I_{\text{DL}}$ is the maximum mutual information achievable in SemCom systems at any SNRs. Finally, the mutual information margin between the upper bound and the proposed SemHARQ can be formulated as
${\color[rgb]{0,0,0}\Gamma}$ $={\color[rgb]{0,0,0}I_{\text{U}}-\sum_{k=1}^{K}\max_{\alpha,\beta,\alpha^{-1},{\beta_{k}^{-1}},{\phi_{k}}}I\left(\bm{s};\bm{c}_{k}\right),}$ (32)
where $I\left(\bm{s};\bm{c}_{k}\right)$ denotes the mutual information between the semantic feature vector $\bm{s}$ at the transmitter and the $k$-th task prediction vector in SemHARQ, $\Gamma$ represents the discrepancy between the proposed SemHARQ and the optimal performance the system can achieve with certain deep learning encoding/decoding functions. To further illustrate the performance margin, Fig. 14 presents $I_{\text{U}}$ and the estimated mutual information in SemHARQ. It shows that at high SNR regime, e.g., at $8$ dB, the mutual information of SemHARQ is only $1.7\%$ lower than $I_{\text{U}}$, indicating the effectiveness of the designed system. While at 6-6 dB, the mutual information of SemHARQ is nearly $50\%$ of the upper bound. The poor performance may be due to the constant-coding rate commonly considered for SemCom systems, although the designed HARQ mechanism can fix the constant-coding rate problem for a certain extent. The performance could be further improved with an SNR-adaptive coding system, which can be an interesting topic for further research and design.